National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 1061–1080 of 2520Page 54 of 100
NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1061

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Consider the following relations for two events E and F:

  1. P(E ∩ F) ≥ P(E) + P(F) - 1
  2. P(E ∪ F) = P(E) + P(F) + P(E ∩ F)
  3. P(E ∪ F) ≤ P(E) + P(F)

Which of the above relations is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    3 only

  3. COption (c)

    1 and 3 only

  4. DOption (d)

    1, 2 and 3

View answer & explanation

Correct answer

Answer: (c)C.

1 and 3 only

Why this answer is correct

Given:

  1. P(E ∩ F) ≥ P(E) + P(F) - 1
  2. P(E ∪ F) = P(E) + P(F) + P(E ∩ F)
  3. P(E ∪ F) ≤ P(E) + P(F)

Calculation:

(1) P(E ∩ F) ≥ P(E) + P(F) - 1

The above inequality is the lower bound of the intersection of two events E and F. This relation is called the Principle of Inclusion-Exclusion (for two events).

If you think of this intuitively, the left-hand side P(E ∩ F) represents the probability of both events E and F happening together. In the worst case, this can't be higher than the sum of the probabilities of E and F occurring individually (i.e., P(E) + P(F)), less the probability of at least one of them happening (which is 1). So, the relation holds true that the intersection should be greater than or equal to the sum of probabilities of two events E and F, minus 1. (2) P(E ∪ F) = P(E) + P(F) + P(E ∩ F)

The correct formula -

P(E ∪ F) = P(E) + P(F) - P(E ∩ F)

The above equation is the formula for the probability of the union of two events E and F. The original relation stated was incorrect because the plus sign before the last term P(E ∩ F) should be a minus sign.

This formula tells us the total probability of either event E, event F, or both happening is equal to the sum of their individual probabilities minus the probability of them both occurring (since we double-counted that in the sum). (3) P(E ∪ F) ≤ P(E) + P(F)

The above inequality is the upper bound on the union of two events E and F. Essentially, this is stating that the probability of either E happening, or F happening, or both, is naturally less than or equal to the sum of their individual probabilities.

If E,F are mutually exclusive (cannot occur at the same time), the sign becomes equality instead of inequality.

So, statement 1 and 3 are correct. Hence option (3) is correct.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1062

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If P(A|B) < P(A), then which one of the following is correct?

  1. AOption (a)

    P(B|A) < P(B)

  2. BOption (b)

    P(B|A) > P(B)

  3. COption (c)

    P(B|A) = P(B)

  4. DOption (d)

    P(B|A) > P(A)

View answer & explanation

Correct answer

Answer: (a)A.

P(B|A) < P(B)

Why this answer is correct

Formula used:

P(A|B) = \(\rm \frac{P(A ∩ B)}{P(B)}\)

P(B|A) = \(\rm \frac{P(A ∩ B)}{P(A)}\)

Explanation:

Given that,

P(A|B) < P(A)

⇒ \(\rm \frac{P(A ∩ B)}{P(B)}\) < P(A)       [∵ P(A|B) = \(\rm \frac{P(A ∩ B)}{P(B)}\)]

⇒ P(A ∩ B) < P(A).P(B)

⇒ \(\rm \frac{P(A ∩ B)}{P(A)}\) < P(B)

⇒ P(B|A) < P(B)       [∵ P(B|A) = \(\rm \frac{P(A ∩ B)}{P(A)}\)]

∴  P(B|A) < P(B)

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1063

National Defence Academy (NDA) 2021• Mathematics• Mathematics

When the measure of central tendency is available in the form of mean, which one of the following is the most reliable and accurate measure of variability?

  1. AOption (a)

    Range

  2. BOption (b)

    Mean deviation

  3. COption (c)

    Standard deviation

  4. DOption (d)

    Quartile deviation

View answer & explanation

Correct answer

Answer: (c)C.

Standard deviation

Why this answer is correct

Dispersion is the extent to which values in a distribution differ from the average of the distribution. To quantify the extent of the variation, there are certain measures namely:

  • Measures of Dispersion From Average:
  • ​The mean deviation (MD) of a series is the arithmetic average of the deviation of various items from a measure of central tendency (mean, median, and mode).
  • Standard Deviation (S. D.): One of the most stable measures of variability, it is the most important and commonly used measure of dispersion. It measures the absolute dispersion or variability of a distribution. Standard deviation is the positive square root of the mean of the squared deviations of observations from their mean. It is denoted by S.D. or σx.
  • Measures Based Upon Spread of Values:
  • Range: It is the simplest measure of variability both from the point of view of calculation and understanding. It is the difference between the largest and the smallest observations in the data. For example, the range of scores 2, 5, 6, 4, 12, 10, 9 and 8 is 12-2 = 10.
  • Interquartile range (IQR) is a measure of variability, based on dividing a data set into quartiles. It divides a rank-ordered data set into four equal parts. Q1 is the mid-value in the first half of the rank-ordered data, Q2 is the score that has exactly two quarters (median). Q3 is the mid-value in the second half of the rank-ordered data. 

For example, consider the number 3, 4, 6, 9, 11, 12, 14, 15.

Q1 is the middle value in the first half of the data set.

The middle value is the average of the two middle values; that is, Q1 = \(4+6\over 2\) = 5.

Q3 is the middle value in the second half of the data set.

i. e. Q3 = \(12+14\over 2\) = 13

Semi Quartile Range = \(Q3 - Q1\over 2\) = \(13 - 5 \over 2\) = 4

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1064

National Defence Academy (NDA) 2021• Mathematics• Mathematics

A problem is given to three students A, B and C, whose probabilities of solving the problem independently are \(\rm \frac{1}{2}\), \(\rm \frac{3}{4}\) and p respectively. If the probability that the problem can be solved is \(\rm \frac{29}{32}\), then what is the value of p?

  1. AOption (a)

    \(\rm \frac{2}{5}\)

  2. BOption (b)

    \(\rm \frac{2}{3}\)

  3. COption (c)

    \(\rm \frac{1}{3}\)

  4. DOption (d)

    \(\rm \frac{1}{4}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\rm \frac{1}{4}\)

Why this answer is correct

Formula used:

Probability: The probability of an event can be calculated by simply

dividing the favorable number of outcomes by the total number

of possible outcomes.

The value of the probability of an event to happen can lie between 0 and 1 i.e. P(̅{E})

P(E)+P(̅{E}) = 1

\(\rm P(E) + P(\overline{E}) = 1\)

The event \(\rm P(\overline{E}) \) means ‘Not E’.

Calculation:

P(problem is solved) = 1 - P(problem is not solved)

⇒ P(problem is solved) = 1 - P(A̅) × P(B̅) P(C̅)

⇒ \(\rm (\frac{29}{32}) = 1-(1 - \frac{1}{2})(1 - \frac{3}{4})(1 - p)\)

⇒ \(\rm (1 - \frac{29}{32}) = (1 - \frac{1}{2})(1 - \frac{3}{4})(1 - p)\)

⇒ \(\rm \frac{3}{32} = \frac{1}{2}× \frac{1}{4} × (1 - p)\)

⇒ 1 - p = \(\rm \frac{3}{4}\)

⇒ P = 1 - \(\rm \frac{3}{4}\)

⇒ P = \(\rm \frac{1}{4}\)  

∴ The value of p is \(\rm \frac{1}{4}\).

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1065

National Defence Academy (NDA) 2021• Mathematics• Mathematics

In a cricket match, a batsman hits a six 8 times out of 60 balls he plays. What is the probability that on a ball played he does not hit a six?

  1. AOption (a)

    \(\rm \frac{2}{3}\)

  2. BOption (b)

    \(\rm \frac{1}{15}\)

  3. COption (c)

    \(\rm \frac{2}{15}\)

  4. DOption (d)

    \(\rm \frac{13}{15}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\rm \frac{13}{15}\)

Why this answer is correct

Formula used:

Probability of occurrence of the event:

P(E) =  \(\frac{n(E)}{n(S)}\)

Where,

n(E) = Number of favorable outcome

n(S) = Number of possible outcome

Calculation:

A batsman hits sixes 8 times out of 60 balls then

Probability of getting sixes

P(A) = \(\rm \frac{8}{60}\) = \(\rm \frac{2}{15}\)​

The probability that he does not hit a six

P(A̅) = 1 − \(\rm \frac{2}{15}\) 

⇒ P(A̅) = \(\rm \frac{13}{15}\)

∴ Required probability is \(\rm \frac{13}{15}\).

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1066

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Direction: Consider the following for the next two (02) items that follow.

Two regression lines are given as 3x - 4y + 8 = 0 and 4x - 3y - 1 = 0.

Consider the following

Statements:

  1. The regression line of y on x is \(\rm y = \frac{3}{4}x+2\)
  2. The regression line of x on y is \(\rm x = \frac{3}{4}y+\frac{1}{4}\)

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (c)C.

Both 1 and 2

Why this answer is correct

Concept:

The regression line of Y on X: 

Y =  b1X + b0 

The regression line of X on Y: 

X = a1Y +  a0 

Where,

b0, a0 is constant and b1, a1 is the regression coefficient.

Calculation:

We have,

3x - 4y + 8 = 0

⇒ 4y = 3x + 8

⇒ \(y = \frac{3}{4}x+2\)

Also,

4x - 3y - 1 = 0

⇒ 4x = 3y + 1

⇒ x = \(\rm \frac{3y}{4} + \frac{1}{4}\)

∴ Both 1 and 2 statements are correct.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1067

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Direction: Consider the following for the next two (02) items that follow.

Two regression lines are given as 3x - 4y + 8 = 0 and 4x - 3y - 1 = 0.

Consider the following

Statements:

  1. The coefficient of correlations r is \(\rm \frac{3}{4}\).
  2. The means of x and y are 3 and 4 respectively.

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (a)A.

1 only

Why this answer is correct

Formula used:

Coefficient of correlations r = \(\rm ± \sqrt{b_{xy}b_{yx}}\)

The regression coefficient bxy & byx are the change occurring in x for a

unit change in y and y for a unit change in x respectively.

Calculation:

3x - 4y + 8 = 0  

⇒ 4y = 3x + 8   

⇒ y = \(\rm \frac{3}{4}\)x + 2     ----(1)

⇒  byx = \(\rm \frac{3}{4}\)

Similarly,

4x - 3y - 1 = 0   -----(2)

bxy = \(\rm \frac{3}{4}\)

We know that,

r = \(\pm\sqrt{\frac{3}{4}\times \frac{3}{4}}\)

⇒ r = \(\rm \frac{3}{4}\)

Hence, statement 1 is correct.

Put the value of y from the equation (1) in (2), 

4x - 3(​\(\frac{3}{4}\)x + 2) - 1 = 0 

⇒ 16x - 9x - 28 = 0

⇒ 7x = 28

⇒ x = 4

Hence, from equation (2),

4(4) - 3y - 1 = 0

⇒ 3y = 15

⇒ y = 5

Hence, statement 2 is incorrect.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1068

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Direction: Consider the following for the next two (02) items that follow.

The marks obtained by 60 students in a certain subject out of 75 are given below:

MarksNumber of
Students
15 - 204
20 - 255
25 - 3011
30 - 356
35 - 405
40 - 458
45 - 509
50 - 556
55 - 604
60 - 652

What is the median?

  1. AOption (a)

    35

  2. BOption (b)

    38

  3. COption (c)

    39

  4. DOption (d)

    40

View answer & explanation

Correct answer

Answer: (c)C.

39

Why this answer is correct

Formula used:

Median of grouped data:

Median = \(\rm l + \frac{\frac{N}{2} - C}{f} × h\)

Where,

l = lower limit of the median class

N = number of observations

f = Frequency of median class

h = Class size

C = Cumulative frequency of class preceding the median class.

To find the median class:  If N is odd, the median is the [(N+1)/2].

If n is even, then the median will be the average of the n/2th and

the (n/2 +1)th observation.

Calculation:

MarksNumber of
Students(fi)
Cumulative Frequency(C.F)
15-2044
20-2559
25-301120
30-35626
35-40531
40-45839
45-50948
50-55654
55-60458
60-65260
∑f = N = 60

 

N = 60, \(\rm \frac{N}{2} = \frac{60}{2} = 30\)

Median class is 35 - 40

l = 35, f = 5, h = 5 and C = 26

⇒ Median = 35 + \(\rm \frac{30 -26}{5} × 5\)

⇒ Median = 35 + 4 = 39

∴ Required median is 39.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1069

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Direction: Consider the following for the next two (02) items that follow.

The marks obtained by 60 students in a certain subject out of 75 are given below:

MarksNumber of
Students
15 - 204
20 - 255
25 - 3011
30 - 356
35 - 405
40 - 458
45 - 509
50 - 556
55 - 604
60 - 652

What is the mode?

  1. AOption (a)

    27.27

  2. BOption (b)

    27.73

  3. COption (c)

    27.93

  4. DOption (d)

    28.27

View answer & explanation

Correct answer

Answer: (b)B.

27.73

Why this answer is correct

Formula used:

Mode = L +  \([ \frac{f_{1} - f_{0}}{2f_{1} - f_{0} - f_2}]\times h\)

Where,

l = lower limit of the modal class

h = size of the class interval

f1 = frequency of the modal class

f0 = frequency of the class preceding the modal class

f2 = frequency of the class succeeding the modal class

Let us take an example to understand this clearly.

 

Calculation:

From the table, we can see

Class with the highest frequency is 25 - 30 = modal class

Hence,

L = 25, f1 = 11, f0 =  5, f2 = 6, h = 5

Therefore,

Mode = 25 + \([\frac{11 - 5}{2\times 11 - 5 - 6}]\times 5\)

⇒ Mode = 27.73

∴ The required mode is 27.73

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1070

National Defence Academy (NDA) 2021• Mathematics• Mathematics

What is the mean of natural numbers in the interval [15, 64]?

  1. AOption (a)

    36.8

  2. BOption (b)

    38.3

  3. COption (c)

    39.5

  4. DOption (d)

    40.3

View answer & explanation

Correct answer

Answer: (c)C.

39.5

Why this answer is correct

Concept:

Sum of n terms of AP with first and last term a & d respectively is

S = \(\frac{n}{2}(a +l)\)

Mean: It is the ratio of the sum of all observations and the total

number of observations. 

Calculation:

The given interval is [15, 64].

Hence, numbers in the interval are

15, 16, ......63, 64 which is AP.

a = 15, l = 64

Let there are n numbers in the series.

⇒ Sum = \(\frac{n}{2}(a +l)\)

⇒ S = \(\frac{n}{2}(15 +64)=\frac{79n}{2}\)

Hence, 

Mean = \(\frac{1}{n}( \frac{79n}{2})\)

⇒ Mean = 39.5

∴ The mean of natural numbers in the interval [15, 64] is 39.5.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1071

National Defence Academy (NDA) 2021• Mathematics• Mathematics

For the set of numbers x, x, x + 2, x + 3, x + 10 where x is a natural number, which of the following is/are correct?

  1. Mean > Mode
  2. Median > Mean

Select the correct answer using the code given below.

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (a)A.

1 only

Why this answer is correct

Formula used:

Mean: It is the ratio of the sum of all observations and the total

number of observations.

Mode: The mode is the value that is repeatedly occurring in a given set.

Relation b/w mean, mode and median:

Mode = 3(Median) - 2(Mean)

Calculation:

Given set of numbers are

x, x, x + 2, x + 3, x + 10 

According to queston, x is a natural number, then put x = 1

Set of numbers are

1, 1, 3, 4, 11

Mean = \(\rm \frac{1 + 1 + 3 + 4 + 11}{5}\) 

Mean = 4

Mode = 1

Mean > Mode

Hence,

Statement 1 is correct.

We know that,

Mode = 3(Median) - 2(Mean)

⇒ 1 = 3(Median) - 2(4)

⇒ 3(Median) = 9

⇒ Median = 3

Median < Mean

∴ Statement 2 is  incorrect.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1072

National Defence Academy (NDA) 2021• Mathematics• Mathematics

The mean of 10 observations is 5.5. If each observation is multiplied by 4 and subtracted from 44, then what is the new mean?

  1. AOption (a)

    20

  2. BOption (b)

    22

  3. COption (c)

    34

  4. DOption (d)

    44

View answer & explanation

Correct answer

Answer: (b)B.

22

Why this answer is correct

Formula used:

Mean: It is the ratio of the sum of all observations and the total

number of observations.

Property of mean:

Let mean of  a1, a2,  ......an  = A, then

  • Mean of  a1 + k,a2,  + k  ......an + k = A + k
  • Mean of  a1 - k, a2  - k  ......an  - k = A - k
  • Mean of  ka1, ka2,  ......kan  = kA
  • Mean of a1/k  a2/k  ......an/k  = A/k

 

Calculation:

Given that, a mean of 10 observations 

 A = 5.5

According to the question,  each observation is multiplied by 4 and

subtracted from 44. Let new mean is A', then using the property of

mean,

⇒ A' = 44 - 4A

⇒ A' = 44 - 4 × 5.5      (∵ A = 5.5)

⇒ A' = 44 - 22  

⇒ A' = 22

∴  The required mean is 22.

Alternate Method

Given that,

Mean of 10 observation = 5.5

Let a1, a2,  ......a10  are 10 observations.

⇒ (a1 + a2 +   ......a10)/10 = 5.5

⇒ (a1 + a2 +   ......a10) = 55     ------(1)

According to the question, 

Each observation is multiplied by 4 and subtracted from 44. Then,

Mean = [10 ×44 - 4(a1 + a2 +   ......a10)]/10

Using equation (1)

Mean = [10 ×44 - 4 × 55]/10

⇒ Mean = (440 - 220)/10 

⇒ Mean = 220/10 = 22    

∴ The new mean is 22.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1073

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If g is the geometric mean of 2, 4, 8, 16, 32, 64, 128, 256, 512, 1024, then which one of the following is correct?

  1. AOption (a)

    8 < g < 16

  2. BOption (b)

    16 < g < 32

  3. COption (c)

    32 < g < 64

  4. DOption (d)

    g > 64

View answer & explanation

Correct answer

Answer: (c)C.

32 < g < 64

Why this answer is correct

Formula used:

For n terms,  a1, a2, a3, ……., an

Geometric mean = \(\rm ({a_{1}.a_{2}.a_{3}....a_{n}})^{\frac{1}{n}}\)

Calculation:

g is the geometric mean of 2, 4, 8, 16, 32, 64, 128, 256, 512, 1024

⇒ g = \((2 × 4 × 8 × 16 × 32 × 64 × 128 × 256 × 512 × 1024)^{\frac{1}{10}}\)  

⇒ g = (2 × 22 × 23 × 24 × 25 × 26 × 27 × 28 × 29 × 210\()^{\frac{1}{10}}\)

⇒ g = \(\rm 2^{5.5}\)

⇒ 25 < g < 26

∴ 32 < g < 64 is correct.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1074

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If the harmonic mean of 60 and x is 48, then what is the value of x?

  1. AOption (a)

    32

  2. BOption (b)

    36

  3. COption (c)

    40

  4. DOption (d)

    44

View answer & explanation

Correct answer

Answer: (c)C.

40

Why this answer is correct

Formula used:

For n terms,  a1, a2, a3, ……., an.

Harmonic mean = \(\rm \frac{n}{\frac{1}{a_{1}} + \frac{1}{a_{2}} + \frac{1}{a_{3}} + ....... \frac{1}{a_{n}} }\)

If b is the Harmonic mean of a and b then 

b = 2ab/(a + b)

Calculation:

Given, The harmonic mean of 60 and x is 48

According to the formula used 

\(\rm \frac{2\times 60x}{x + 60} = 48\)

⇒ 120x = 48(x + 60)

⇒120x = 48x + 2880

⇒ 120x - 48x = 2880

⇒ x = 40

∴ The value of x is 40.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1075

National Defence Academy (NDA) 2021• Mathematics• Mathematics

What is the mean deviation of first 10 even natural numbers?

  1. AOption (a)

    5

  2. BOption (b)

    5.5

  3. COption (c)

    10

  4. DOption (d)

    10.5

View answer & explanation

Correct answer

Answer: (a)A.

5

Why this answer is correct

Concept:

Mean deviation = \(\frac{1}{n}|x_i-\bar{X}|\)

Where,

\(\bar{X}= \frac{∑ x}{n}\) = mean, 

∑x = sum of all observations,

n = number of observations. 

Calculation:

First 10 even natural numbers are

2, 4, 6, 8, 10, 12, 14, 16, 18, 20

Mean, 

X̅  = \(\rm \frac{2 + 4 + 6 + 8 + 10 + 12 + 14 + 16 + 18 + 20}{10}\) 

⇒ X̅ = 11

Therefore, the mean value is 11

Now, subtract each mean from the first 10 even natural numbers, and ignore the minus symbol if any

⇒ Mean deviation = \(\frac{1}{n}|x_i-\bar{X}|\)

⇒ Mean deviation \(=\frac{1}{10}(|2 - 11|+ |4 - 11| + |6-11|+|8-11|+|10-11|+\|12-11|+|14-11|+|16-11|+|18-11|+|20-11|\)

⇒  \(\rm\frac{9 + 7 + 5 + 3 + 1 + 1 + 3 + 5 + 7 + 9}{10}\) = \(\frac{50}{10}\)

∴ The mean deviation of the first 10 even natural numbers is 5.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1076

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If \(\rm \displaystyle \sum_{i = 1}^{10}x_i = 110\) and \(\rm \displaystyle \sum_{i = 1}^{10}x_i^2 = 1540\)  then what is the variance?

  1. AOption (a)

    22

  2. BOption (b)

    33

  3. COption (c)

    44

  4. DOption (d)

    55

View answer & explanation

Correct answer

Answer: (b)B.

33

Why this answer is correct

Given:

\(\rm \displaystyle \sum_{i = 1}^{10}x_i = 110\)

\(\rm \displaystyle \sum_{i = 1}^{10}x_i^2 = 1540\)

Formula used:

Variance:

\(V(X)=\rm \frac{ \displaystyle \sum_{i = 1}^{n}x_i^2}{n} -\frac{ (\displaystyle \sum_{i = 1}^{n}x_i)^{2}}{n^2} \)

Calculation:

We know that, variance,

\(V(X)=\rm \frac{ \displaystyle \sum_{i = 1}^{n}x_i^2}{n} -\frac{ (\displaystyle \sum_{i = 1}^{n}x_i)^{2}}{n^2} \)

⇒ V(X) = \(\rm \frac{1540}{10} - (\frac{110}{10})^2\)

⇒ V(X) = 154 - 121

⇒ V(X) = 33

∴ The variance is 33.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1077

National Defence Academy (NDA) 2021• Mathematics• Mathematics

3-digit numbers are formed using the digits 1, 3, 7 without repetition of digits. A number is randomly selected. What is the probability that the number is divisible by 3?

  1. AOption (a)

    0

  2. BOption (b)

    \(\rm \frac{1}{3}\)

  3. COption (c)

    \(\rm \frac{1}{4}\)

  4. DOption (d)

    \(\rm \frac{1}{8}\)

View answer & explanation

Correct answer

Answer: (a)A.

0

Why this answer is correct

Concept:

  • Divisibility rule of 3 : A number is completely divisible by 3 if the sum of its digits is divisible by 3
  • nPr = n!/(n - r)!

Calculation:

Total number of three-digit numbers = 3P2 = 3!/(3 - 2)! = 6

Total number of three-digit numbers are 137, 173, 371, 317, 713, 731 

Total number of outcomes = 6 

Since Sum of digits is 11 which is not a multiple of 3 

So, The favorable outcome is 0

**∴ The probability that the number is divisible by 3 is 0/6 = 0.**​

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1078

National Defence Academy (NDA) 2021• Mathematics• Mathematics

What is the probability that the roots of the equation x2 + x + n = 0 are real, where n ∈ ℕ and n < 4?

  1. AOption (a)

    0

  2. BOption (b)

    \(\rm \frac{1}{4}\)

  3. COption (c)

    \(\rm \frac{1}{3}\)

  4. DOption (d)

    \(\rm \frac{1}{2}\)

View answer & explanation

Correct answer

Answer: (a)A.

0

Why this answer is correct

Concept:

Consider the quadratic equation ax2 + bx + c = 0

Discriminant (D): D = b2 - 4ac

For real root: D ≥ 0

b2 - 4ac ≥ 0

Calculation:

Given equation 

x2 + x + n = 0

Finding discriminant: (1)2 - 4n ≥ 0

1  ≥  4n

n ≤ \(\rm \frac{1}{4}\)

There is no value which belongs in 1, 2, 3, ..... (i.e. natural number)

This is possible for n = 0

∴ The probability of the required event is zero.

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1079

National Defence Academy (NDA) 2021• Mathematics• Mathematics

If A and B are two events such that P(not A) = \(\rm \frac{7}{10}\), P(not B) = \(\rm \frac{3}{10}\) and P(A|B) = \(\rm \frac{3}{14}\), then what is P(B|A) equal to?

  1. AOption (a)

    \(\rm \frac{11}{14}\)

  2. BOption (b)

    \(\rm \frac{9}{14}\)

  3. COption (c)

    \(\rm \frac{1}{4}\)

  4. DOption (d)

    \(\rm \frac{1}{2}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\rm \frac{1}{2}\)

Why this answer is correct

Given:

P(not A) = \(\rm \frac{7}{10}\), and P(not B) = \(\rm \frac{3}{10}\)

P(A|B) = \(\rm \frac{3}{14}\)

Formula used:

  • \(\rm P(B|A) = \frac{P(A\cap B)}{P(A)}\)
  • \(\rm P(A|B) = \frac{P(A\cap B)}{P(B)}\)
  • P(A) = 1 - P(not A)
  • P(B) = 1 - P(not B)

 

Calculation:

We have,

\(\rm P(\bar{B}) = 0.3, P(\bar{A}) = 0.7\)

⇒ P(B) = 1 - 0.3 = 0.7 and

P(A) = 1 - 0.7 = 0.3

⇒ P(A) = 0.3      ----(1)

We know that,

\(\rm P(A|B) = \frac{P(A\cap B)}{P(B)}\)

⇒ \(\frac{3}{14} \) = \(\frac{P(A\cap B)}{0.7}\)       (∵ P(A|B) = \(\rm \frac{3}{14}\))

⇒ \(\rm P(A\cap B)\) = 0.15      ----(2)   

We know that,

\(\rm P(B|A) = \frac{P(A\cap B)}{P(A)}\)

⇒ P(B|A) = \(\rm \frac{0.15}{0.3}\)        [From equation (1)& (2)]

⇒ P(B|A) = 0.5 = \(\rm \frac{1}{2}\)

∴  P(B|A) is equal to \(\rm \frac{1}{2}\)

NDA Mathematics 14 Nov 2021 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1080

National Defence Academy (NDA) 2021• Mathematics• Mathematics

Seven white balls and three black balls are randomly placed in a row. What is the probability that no two black balls are placed adjacently?

  1. AOption (a)

    \(\rm \frac{7}{15}\)

  2. BOption (b)

    \(\rm \frac{8}{15}\)

  3. COption (c)

    \(\rm \frac{11}{15}\)

  4. DOption (d)

    \(\rm \frac{13}{15}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\rm \frac{7}{15}\)

Why this answer is correct

Concept:

Probability:

  • The probability of an event can be calculated by simply dividing the favorable number of outcomes by the total number of possible outcomes.
  • The value of the probability of an event to happen can lie between 0 and 1.

Combination Formula:

 nCr = \(\rm \frac{n!}{r!(n-r)!}\)

nCr = number of combinations

n = total number of objects in the set

r =number of choosing objects from the set

Calculation:

The number of ways of arranging 10 balls in a row = 10 P10.​=10!

Now, the condition is to not arrange two black balls adjacently.

So, first, arrange the 7 while balls first and arrange these 3 black balls in between the created 8 places between each white ball.

So, number of favorable outcomes of arranging 3 black balls and 7 white balls

= 8 P3 × 7!

Thus, the probability that no two black balls are placed adjacently is

P = (8 P3 × 7!)​ / 10 P10

⇒ P = \(\rm \frac{7}{15}\)

∴ The required probability is \(\rm \frac{7}{15}\).

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.